A single-phase asymmetrical semi-converter feeds an R-L load with R=50 Ω and a large L= 10 mH so that load current is constant. The source voltage is 230 V, 50 Hz. For a firing angle of 600, determine: (i) Average value of output voltage and output current. (ii) Average and rms values of diode, thyristor and source currents.(iii) Input power factor. 


1
Expert's answer
2021-04-26T02:03:54-0400

i)Average value of output voltage and output current.

"V_O=\frac{v_m}{\pi}=\frac{230 \sqrt{2}}{\pi}(1+cos 60)=155.31V"

"I_o=\frac{V_o}{R}=\frac{155.31}{50}=3.11A"


(ii) Average and rms values of diode, thyristor and source currents.

"I_pavg=3.11\times\frac{\pi-\frac{\pi}{3}}{2\pi}=1.04A"

"I_DAvg=3.11\times\frac{\pi+\frac{\pi}{3}}{2\pi}=2.07A"

"I_Trms=3.11\times\sqrt{\frac{\pi-\pi /3}{2\pi}}=1.79A"

"I_Drms=3.11\times\sqrt{\frac{\pi+\pi /3}{2\pi}}=2.54A"

Source current "I_DAvg=3.11\times\frac{\pi+\frac{\pi}{3}}{2\pi}=2.07A"

"I_Drms=3.11\times\sqrt{\frac{\pi+\pi /3}{2\pi}}=2.54A"

(iii) Input power factor. 

"\frac{I_{SI}}{I_{Sr}}cos\frac{\alpha}{2}=\frac{2.42}{2.54}cos\frac{60}{2}=0.825"